OpenStudy (anonymous):

Can someone check my solution? Question: 1/2(e^2)(y)= x^3+2; solve for Y in terms of X. What I got: y=ln(x^3+2)

OpenStudy (anonymous):

y = 2*(x^3+2)/(e^2)

ganeshie8 (ganeshie8):

is the equation like below : \(\huge \frac{1}{2}e^{2y} = x^3+2\) ?

ganeshie8 (ganeshie8):

@hitech89

OpenStudy (anonymous):

ahh. if so, the answer is y=(1/2)ln(2x^3+4)